OpenStudy (anonymous):

What is the value of the expression 1/3^-5 ? A. 243 B. 15 C.1/15 D.1/243

OpenStudy (anonymous):

@aaronq

Nnesha (nnesha):

\[\huge\rm x^{-m} = \frac{ 1 }{ x^m } or \frac{ 1 }{ x^{-m} } = x^m\] when there is a negative exponent you need to flip the base to the other side

OpenStudy (anonymous):

so witch one is it

Nnesha (nnesha):

that what i should ask so which one is it ?

Nnesha (nnesha):

s*

OpenStudy (anonymous):

d

Nnesha (nnesha):

you have to flip the base in order to convert negative exponent to positive

OpenStudy (anonymous):

so a

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

thank you i was confused